# Multi-modal Misinformation Detection: Approaches, Challenges and Opportunities

### v6

@@ 3.3.4. Graph neural network architectures
In recent years, Graph Neural Networks (GNNs) have been successfully exploited for fake news detection  (Cui et al., 2020; Song et al., 2021c; Benamira et al., 2019), thereby they have caught researchers’ attention for multi-modal misinformation detection tasks as well. In this category of deep learning solutions, article content e.g., text, image etc. are represented by graphs and then graph neural networks are used to extract the semantic-level features.
For instance, Wang et al. construct a graph for each social media post based on the point-wise mutual information (PMI) score of pairs of words, extracted objects in visual content and knowledge concept through knowledge distillation. Then, utilize a Knowledge-driven Multi-modal Graph Convolutional Network (KMGCN) which extracts the multi-modal representation of each post through Graph convolutional networks (Wang et al., 2020).
Another GCN based model is GAME-ON (Dhawan et al., 2022) which represents each news with uni-modal visual and textual graphs and then projects them into a common space. To capture multi-modal representations, Game-on applies a graph attention layer on a multi-modal graph generated out of modality graphs.
@@ 3.3.5. Cross-modal discordance-aware architectures
In the previous categories we discussed above, deep learning models are used to fuse different modalities in order to obtain discriminating representations. However, in this category, deep learning architectures are customized based on some identified discordance between modalities. The intuition is that: fabrication of either modality will lead to dissonance between the modalities and results in misrepresented, misinterpreted and misleading news. Therefore, there are subtle cross-modal discordance clues that could be identified and learned by customized architectures. In light of this, methods that employ “contrastive learning” or Contrastive Language-Image Pre-Training (CLIP) based architectures may be included in this category (Cui et al., 2023; Jiang et al., 2023).
In many cases, fake news propagators use irrelevant modalities e.g., image, video, audio etc., for false statements to attract readers’ attention. Thus, the similarity of text to other modalities e.g., image, audio etc., is a cue for measuring the credibility of a news article.
With that said, Zhou et al. (Zhou et al., 2020b), propose SAFE, a Similarity-Aware Multi-Modal Fake News Detection framework by defining the relevance between news textual and visual information using a modified cosine similarity.
Similarly, Giachanou et al., propose a multi-image system that combines textual, visual and semantic information (Giachanou et al., 2020). The semantic representation refers to the text-image similarity calculated using the cosine similarity between the title and image tag embeddings.
In another work, Singhal et al. (Singhal et al., 2021) develop an inter-modality discordance based fake news detector which learns discriminating features and employs a modified version of contrastive loss that explores the inter-modality discordance. Xue et al. (Xue et al., 2021), propose a Multi-modal Consistency Neural Network (MCNN) which utilizes a similarity measurement module that measures the similarity of multi-modal data to detect the possible mismatches between the image and text. Lastly, Biamby et al.  (Biamby et al., 2021), leverage CLIPmodel (Radford et al., 2021), to jointly learn image/text representation to detect Image-Text inconsistencies in Tweets. Instead of concatenating vector representation, CLIP jointly trains an image encoder and a text encoder to predict the correct pairings of a batch of (image, text) training examples.
@@ 3.3.6. Foundation models and prompt-based techniques
A foundation model is a large machine learning model that is trained on large scale datasets such that it can be adapted to a wide range of downstream tasks. Some examples of multi-modal foundation models are pre-trained GPT-4 (OpenAI, 2023), DALL-E (Ramesh et al., 2021), Florence (Yuan et al., 2021), Flamingo (Alayrac et al., 2022) and so on. In-Context Learning (ICL) is the simplest and one of the most effective ways of using foundation models. ICL is a training-free technique where models learn to learn from limited demonstrations and descriptions and generalize to unseen tasks (Tai et al., 2023). The learn to learn concept was first introduced in Meta-learning which is a family of machine learning techniques that uses few examples to adapt the model to new tasks. In recent years meta-learning has been used for different applications including multi-modal misinformation detection (Wang et al., 2021a; Yue et al., 2023). However, GPT-3 paper (Brown et al., 2020) shows that few-shot learning is an emergent capability of Large Language Models (LLMs) and could be taken advantage of using ICL techniques. In fact, a frozen model can be conditioned to perform a variety of tasks through ICL, where a user primes the model for a given task through prompt design, i.e., manually crafting a text prompt with descriptions or examples of the task.
@@
A more efficient and effective method for conditioning frozen models is using tunable prompts. Unlike model fine-tuning, which adjusts the model’s parameters by further training on new data, prompt-tuning tunes parameters of the prompt tokens, while keeping the pre-trained model frozen. (Lester et al., 2021).
@@
ICL techniques including few-shot and zero shot prompting, as well as prompt tuning have been widely used to query Large Language Models (LLMs) for a variety of downstream tasks including misinformation detection. For example, Jiang et al. (Jiang et al., 2022) study the role of prompt learning in detecting fake news. In another work (Gao et al., 2023), Gao et al. put forward a prompt-tuning template to extract knowledge from a pretrained LM for detecting misinformation. Another example is a work by Tian et al. (Tian et al., 2023), where few-shot learning is leveraged for troll detection. In another work by Lin et al.  (Lin et al., 2023), prompt tuning is used for rumor detection using a zero-shot framework. Similarly,  (Zuo et al., 2022) presents a continual learning framework that applies prompt tuning for rumor detection.
However, there are few existing works that utilize them for misinformation detection in multi-modal settings. One of the existing works is PromptHate (Cao et al., 2023b), a simple prompt-based multi-modal model, that prompts pre-trained language models (PLMs) for hateful meme classification. PromptHate construct simple prompts and provide a few in-context examples to exploit the implicit knowledge in the pre-trained RoBERTa to classify hateful memes. In another work (Cui et al., 2023), a novel propaganda detection model a.k.a. Antipersuasion Prompt Enhanced Contrastive Learning (APCL) is proposed which is leveraged for detecting propaganda. Prompt is designed with a persuasion prompt template and an anti-persuasion prompt template to build matched text-image and mismatched text-image pairs respectively. Later on, the distances between the two prompt templates and pair of text and image are used for detection. More recently, Cao et al. leverage a pre-trained vision-language models (PVLMs) in a zero-shot and fine-tuning free VQA setting to address the problem of meme detection by generating hateful content-centric image captions (Cao et al., 2023a).
@@
In addition, Jian et al. (Jiang et al., 2023) propose a Similarity-Aware Multimodal Prompt Learning (SAMPLE) framework that incorporates prompt-tuning into multimodal fake news detection. SAMPLE uses three prompt templates: discrete prompting, continuous prompting and mixed prompting to the original input text, and employs the pre-trained RoBERTa to extract text features from the prompt. Furthermore, the pre-trained CLIP is used to obtain the input texts, input images, and their semantic similarities. SAMPLE introduces a similarity-aware multimodal feature fusing approach that applies standardization and a Sigmoid function to adjust the intensity of the final cross-modal representation and mitigate noise injection via uncorrelated cross-modal features.
A summary of the aforementioned deep learning based works is demonstrated in Table 1. It is worth mentioning that many of the state-of-the-art solutions utilize a hybrid of deep learning solutions.
Table 1. A summary of the existing deep learning solutions.
@@
(Alonso-Bartolome and Segura-Bedmar, 2021) Concatenation
@@
(Wang et al., 2021c) Attention Mech.
@@
(Wang et al., 2022) Attention Mech.
@@
(Biamby et al., 2021) Cross-Modal Cue
@@
Multimodal Misinformation Detection Classic Machine Learning Deep Learning Embeddings Concatenation Algebraic Joint Structure Ensemble Learning & Voting Non-convex Optimization Concatenation-based Networks Attention-based Networks Graph Neural Networks Cross-modal Discordance Aware Networks Foundation Models & Prompt-based Techniques
Figure 3. An overview of the multi-modal solutions.
4. Multi-modal data study
Data acquisition and preparation is the most important building block of a machine learning pipeline. Machine learning models leverage training data to continuously improve themselves over time. Thus, sufficient good quality, and in most cases annotated data is extremely crucial for these models to operate effectively. With that said, in this section we introduce and compare some of the existing multi-modal datasets for fake news detection task. Later on we will discuss some of the limitations of these datasets.
Image-Verification-Corpus: [^9] https://githubhelp.com/MKLab-ITI/image-verification-corpus is an evolving dataset containing 17,806 fake and real posts with images shared on Twitter. This dataset is created as an open corpus of tweets containing images that may be used for assessing online image verification approaches (based on tweet texts and user features), as well as building classifiers for new content. Fake and real images of this dataset have been annotated by online sources that evaluate the credibility of the images and the events they are associated with (Boididou et al., 2018).
@@ 4. Multi-modal data study
Fakeddit: [^10] https://github.com/entitize/Fakeddit is a dataset collected from Reddit, a social news and discussion website where users can post submissions on various subreddits. Fakeddit consists of over 1 million submissions from 22 different subreddits spanning over a decade with the earliest submission being from 3/19/2008, and the most recent submission being from 10/24/2019. These subreddits are posted on highly active and popular pages by over 300,000 users. Fakeddit consists of submission titles, images, user comments, and submission metadata including score, the username of the author, subreddit source, sourced domain, number of comments, and up-vote to down-vote ratio. Approximately 64% of the samples have both text and image data (Nakamura et al., 2020). Samples of this dataset are annotated with 2-way, 3-way and 6-way labels including true, satire/parody, misleading content, manipulated content, false connection and imposter content.
NewsBag: comprises 200,000 real news and 15,000 fake articles. The real training articles have been collected from the Wall Street Journal and the fake ones from The Onion website111111 https://www.theonion.com/which publishes satirical content. However, the samples of the test set are collected from different websites i.e., TheRealNews [^12] https://therealnews.com/ and ThePoke [^13] https://www.thepoke.co.uk/. The rational behind using different sources of news for the training and test sets is to observe how well the models could be generalized to unseen data samples. The NewsBag dataset is a highly imbalanced dataset. Thus, to tackle this issue, NewsBag ++ is also released which is the augmented training version of the NewsBag dataset and contains 200,000 real and 389,000 fake news articles. Another weakness of NewsBag dataset is that it does not have any social context information such as spreader information, sharing trends and reactions such as user comments and engagements (Jindal et al., 2020).
$M$-COVID:  [^14] https://github.com/bigheiniu/$M$-COVID is a multi-lingual and multi-dimensional COVID-19 fake news data repository. This dataset comprises 3981 fake news and 7192 trustworthy information in 6 different languages i.e., English, Spanish, Portuguese, Hindi, French and Italian. MM-COVID consists of visual, textual and social context information e.g., users and networks information (Li et al., 2020a). Samples of this dataset are annotated by Snopes151515www.snopes.com and Poynter161616www.poynter.org/coronavirusfactsalliance/ crowed source domains where the experts and journalists evaluate and fact check news content and annotate contents as either fake or real. While Snopes is an independent publication which mainly contains English content, Poynter is an international fact-checking network (IFCN) which unites 96 different fact-checking agencies such as PolitiFact [^17] https://www.politifact.com/ in 40 languages.
ReCOVery: [^18] https://github.com/apurvamulay/ReCOVery contains 2,029 news articles that have been shared on social media, most of which (2,017 samples) have both textual and visual information for multi-modal studies. ReCOVery is imbalanced in news class i.e., the proportion of real vs. fake articles is around 2:1. The number of users who spread real news (78,659) and users sharing fake articles (17,323) is greater than the total number of users included in the dataset (93,761). In this dataset, the assumption is that users can engage in spreading both real and fake news articles. Samples of this dataset are annotated by two factchecking resources: NewsGuard [^19] https://www.newsguardtech.com/ and Media Bias/Fact Check (MBFC) [^20] https://mediabiasfactcheck.com/ which is a website that evaluates factual accuracy and political bias of news media. MBFC labels each news media as one of six factual-accuracy levels based on the fact-checking results of the previously published news articles. Samples of ReCOVery are collected from 60 news domains, from which 22 are the sources of reliable news articles (e.g., National Public Radio [^21] https://www.npr.org/ and Reuters [^22] https://www.reuters.com) and the remaining 38 are sources to collect unreliable news articles (e.g., Human Are Free [^23] http://humansarefree.com/ and Natural News242424www.https://www.naturalnews.com/ (Zhou et al., 2020a).
CoAID:  [^25] https://github.com/cuilimeng/CoAID Covid-19 heAlthcare mIsinformation Dataset or CoAID is a diverse COVID-19 healthcare misinformation dataset, including fake news on websites and social platforms, along with users’ social engagement about the news. It includes 5,216 news, 296,752 related user engagements, 926 social platform posts about COVID-19, and ground truth labels. The publishing dates of the collected information range from December 1, 2019 to September 1, 2020. In total, 204 fake news articles, 3,565 true news articles, 28 fake claims and 454 true claims are collected. Real news articles are crawled from 9 reliable media outlets that have been cross-checked as reliable e.g., National Institutes of Health (NIH)2626262https://www.nih.gov/news-events/news-releases and CDC [^27] https://www.cdc.gov/coronavirus/2019-ncov/whats-new-all.html Fake news is retrieved from several fact-checking websites, such as PolitiFact, Health Feedback [^28] https://healthfeedback.org/ etc  (Cui and Lee, 2020).
Figure 4. Number of news articles by dataset.
MMCoVaR: is a Multi-modal COVID-19 Vaccine Focused Data Repository (MMCoVaR). Articles of this dataset are annotated using two news website source checking and the tweets are fact checked based on stance detection approach. MMCoVaR comprises 2,593 articles issued by 80 publishers and shared between 02/16/2020 05/08/2021 and 24,184 Twitter posts collected between 04/17/2021 to 05/08/2021. Samples of this dataset are annotated by Media Bias Chart and Media Bias/Fact Check (MBFC) and classified into two levels of credibility: reliable and unreliable. Thus, articles are labeled as either credible or unreliable and tweets are annotated as reliable, inconclusive or unreliable (Chen et al., 2021). It is worth mentioning that textual, visual and social context information are available for the news articles.
N24News:  [^29] https://github.com/billywzh717/N24News is a multi-modal dataset extracted from the New York Times articles published from 2010 to 2020. Each news belongs to one of 24 different categories e.g., science, arts etc. The dataset comprises up to 3,000 samples of real news for each category. In total, 60,000 news articles are collected. Each article sample contains a category tag, headline, abstract, article body, image, and corresponding image caption. This dataset is randomly split into training/validation/testing sets in the ratio of 8:1:1. (Wang et al., 2021b). The main weakness of this dataset is that it does not have any fake samples and all of the real samples are collected from a single source i.e., The New York Times.
@@
Table 2. A comprehensive evaluation of classification performance on the Fakeddit dataset focused on accuracy. The study explored different image-text embedders. (Nakamura et al., 2020)
@@
MuMiN:  [^30] https://github.com/MuMiN-dataset/mumin-build Large-Scale Multilingual Multi-modal Fact-Checked Misinformation Social Network Dataset (MuMin) comprises 21 million tweets belonging to 26 thousand Twitter threads, each of which has been linked to 13 thousand fact-checked claims in 41 different languages. MuMiN is available in three large, medium and small versions with largest one consisting of 10,920 articles and 6,573 images. In this dataset, if the claim is “mostly true”, it is labeled as factual. When the claim is deemed “half true” or “half false” it is labeled as misinformation, with the justification that a statement containing a significant part of false information should be considered as a misleading content. When there is no clear verdict then the verdict is labelled as other.  (Nielsen and McConville, 2022).
@@
A summary and a side by side comparison of the aforementioned datasets are illustrated in Fig. 4 and Table 3. As demonstrated, most of these datasets are small in size, are annotated with binary labels, extracted from limited resources such as Twitter, and only contain limited modalities i.e., text and image. Additionally, Table 2 showcases a comparison and evaluation of various methods’ performance on the Fakeddit dataset (Nakamura et al., 2020) [^31] The table is based on the work in (Nakamura et al., 2020) . .
@@
N24News (Wang et al., 2021b) 60,000 24 image,text New York Times All samples are real from 24 different categories
@@
5. Challenges and opportunities in multi-modal misinformation detection
Recent studies on multi-modal learning have made significant contributions to the field of multi-modal fake news detection. However, there are still weaknesses and shortcomings which recognizing them opens the door to new opportunities not only in fake news detection study, but also in the multi-modal field in general. In this section, we discuss challenges, shortcomings and opportunities in multi-modal fake news detection. We provide non-exhaustive lists of challenges and shortcomings for each direction of multi-modal misinformation research.
5.1. Data related challenges

### v7

@@ 3.4.4. Graph Neural Network Architectures
In recent years, Graph Neural Networks (GNNs) have been successfully exploited for fake news detection (Cui et al., 2020; Song et al., 2021c; Benamira et al., 2019), thereby catching researchers’ attention for multi-modal misinformation detection tasks as well. In this category of deep learning solutions, article content (e.g., text, image, etc.) is represented by graphs, and then graph neural networks are used to extract the semantic-level features.
For instance, Wang et al. construct a graph for each social media post based on the point-wise mutual information (PMI) score of pairs of words, extracted objects in visual content, and knowledge concepts through knowledge distillation. They then utilize a Knowledge-driven Multi-modal Graph Convolutional Network (KMGCN) which extracts the multi-modal representation of each post through graph convolutional networks (Wang et al., 2020).
Another GCN-based model is GAME-ON (Dhawan et al., 2022), which represents each news item with uni-modal visual and textual graphs and then projects them into a common space. To capture multi-modal representations, GAME-ON applies a graph attention layer on a multi-modal graph generated out of modality graphs.
@@ 3.4.5. Cross-modal Discordance-aware Architectures
In the previously discussed categories, deep learning models are employed to merge different modalities to create distinguishing representations. However, in this category, deep learning architectures are tailored to address identified discrepancies between modalities. The idea is that fabricating either modality will cause dissonance between them, leading to misrepresented, misinterpreted, and misleading news. Therefore, subtle cross-modal discordance clues can be identified and learned by customized architectures. Consequently, methods utilizing “contrastive learning” or Contrastive Language-Image Pre-Training (CLIP) based architectures (Cui et al., 2023; Jiang et al., 2023) may fall into this category.
In many cases, fake news propagators use irrelevant modalities (e.g., image, video, audio, etc.) for false statements to attract readers’ attention. Thus, the similarity of text to other modalities (e.g., image, audio, etc.) is a cue for measuring the credibility of a news article.
With that said, Zhou et al. (Zhou et al., 2020b) propose SAFE, a Similarity-Aware Multi-Modal Fake News Detection framework by defining the relevance between news textual and visual information using a modified cosine similarity.
Similarly, Giachanou et al. propose a multi-image system that combines textual, visual, and semantic information (Giachanou et al., 2020). The semantic representation refers to the text-image similarity calculated using the cosine similarity between the title and image tag embeddings.
In another work, Singhal et al. (Singhal et al., 2021) develop an inter-modality discordance-based fake news detector which learns discriminating features and employs a modified version of contrastive loss that explores the inter-modality discordance.
Xue et al. (Xue et al., 2021) propose a Multi-modal Consistency Neural Network (MCNN) which utilizes a similarity measurement module that measures the similarity of multi-modal data to detect the possible mismatches between the image and text. Lastly, Biamby et al. (Biamby et al., 2022) leverage the CLIP model (Radford et al., 2021) to jointly learn image/text representation to detect image-text inconsistencies in tweets. Instead of concatenating vector representations, CLIP jointly trains an image encoder and a text encoder to predict the correct pairings of a batch of (image, text) training examples.
@@ 3.4.6. Foundation Models and Prompt-based Techniques
A foundation model is a large machine learning model that is trained on large-scale datasets such that it can be adapted to a wide range of downstream tasks. Some examples of multi-modal foundation models are pre-trained GPT-4 (OpenAI, 2023), DALL-E (Ramesh et al., 2021), Florence (Yuan et al., 2021), Flamingo (Alayrac et al., 2022), and so on.
@@
In-Context Learning (ICL) is the simplest and one of the most effective ways of using foundation models. ICL is a training-free technique where models learn to learn from limited demonstrations and descriptions and generalize to unseen tasks (Tai et al., 2023). The learn-to-learn concept was first introduced in meta-learning, which is a family of machine learning techniques that uses few examples to adapt the model to new tasks. In recent years, meta-learning has been used for different applications, including multi-modal misinformation detection (Wang et al., 2021a; Yue et al., 2023). However, the GPT-3 paper (Brown et al., 2020) shows that few-shot learning is an emergent capability of Large Language Models (LLMs) and could be taken advantage of using ICL techniques. In fact, a frozen model can be conditioned to perform a variety of tasks through ICL, where a user primes the model for a given task through prompt design, i.e., manually crafting a text prompt with descriptions or examples of the task.
@@
A more effective way to condition frozen models is by using tunable prompts. Unlike model fine-tuning, which modifies the model’s parameters through additional training on new data, prompt-tuning adjusts the parameters of the prompt tokens while keeping the pre-trained model frozen (Lester et al., 2021).
@@
ICL techniques, including few-shot and zero-shot prompting, as well as prompt tuning, have been widely used to query Large Language Models (LLMs) for a variety of downstream tasks, including misinformation detection. For example, Jiang et al. (Jiang et al., 2022) study the role of prompt learning in detecting fake news. In another work (Gao et al., 2023), Gao et al. put forward a prompt-tuning template to extract knowledge from a pretrained LM for detecting misinformation. Another example is a work by Tian et al. (Tian et al., 2023), where few-shot learning is leveraged for troll detection. In another work by Lin et al. (Lin et al., 2023), prompt tuning is used for rumor detection using a zero-shot framework. Similarly,  (Zuo et al., 2022) presents a continual learning framework that applies prompt tuning for rumor detection.
However, there are few existing works that utilize them for misinformation detection in multi-modal settings. One of the existing works is PromptHate (Cao et al., 2023b), a simple prompt-based multi-modal model that prompts pre-trained language models (PLMs) for hateful meme classification. PromptHate constructs simple prompts and provides a few in-context examples to exploit the implicit knowledge in the pre-trained RoBERTa to classify hateful memes.
In another work (Cui et al., 2023), a novel propaganda detection model, Antipersuasion Prompt Enhanced Contrastive Learning (APCL), is proposed for detecting propaganda. The prompt is designed with a persuasion prompt template and an anti-persuasion prompt template to build matched text-image and mismatched text-image pairs, respectively. Later on, the distances between the two prompt templates and pairs of text and image are used for detection.
More recently, Cao et al. leverage pre-trained vision-language models (PVLMs) in a zero-shot and fine-tuning-free VQA setting to address the problem of meme detection by generating hateful content-centric image captions (Cao et al., 2023a).
@@
In addition, Jian et al. propose a Similarity-Aware Multimodal Prompt Learning (SAMPLE) framework that incorporates prompt-tuning into multi-modal fake news detection (Jiang et al., 2023). SAMPLE uses three prompt templates: discrete prompting, continuous prompting, and mixed prompting to the original input text, and employs the pre-trained RoBERTa to extract text features from the prompt. Furthermore, the pre-trained CLIP is used to obtain the input texts, input images, and their semantic similarities. SAMPLE introduces a similarity-aware multi-modal feature fusing approach that applies standardization and a Sigmoid function to adjust the intensity of the final cross-modal representation and mitigate noise injection via uncorrelated cross-modal features.
4. Multi-modal Data Study
Data acquisition and preparation are the most important building blocks of a machine learning pipeline. Machine learning models leverage training data to continuously improve themselves over time. Thus, sufficient good quality, and in most cases annotated data, is extremely crucial for these models to operate effectively. With that said, in this section, we introduce and compare some of the existing multi-modal datasets for the fake news detection task. Later on, we will discuss some of the limitations of these datasets.
Image-Verification-Corpus: [^13] https://githubhelp.com/MKLab-ITI/image-verification-corpus is an evolving dataset containing 17,806 fake and real posts with images shared on Twitter. This dataset is created as an open corpus of tweets containing images that may be used for assessing online image verification approaches (based on tweet texts and user features), as well as building classifiers for new content. Fake and real images in this dataset have been annotated by online sources that evaluate the credibility of the images and the events they are associated with (Boididou et al., 2018).
Table 1. A summary of the existing deep learning-based solutions.
@@ 4. Multi-modal Data Study
(Segura-Bedmar and Alonso-Bartolome, 2022) Concatenation
@@
(Wang et al., 2021b) Attention Mech.
@@
(Wang et al., 2022a) Attention Mech.
@@
(Biamby et al., 2022) Cross-Modal Cue
@@
Fakeddit: [^14] https://github.com/entitize/Fakeddit is a dataset collected from Reddit, a social news and discussion website where users can post submissions on various subreddits. Fakeddit consists of over 1 million submissions from 22 different subreddits spanning over a decade, with the earliest submission being from 3/19/2008 and the most recent submission being from 10/24/2019. These subreddits are posted on highly active and popular pages by over 300,000 users. Fakeddit consists of submission titles, images, user comments, and submission metadata including score, the username of the author, subreddit source, sourced domain, number of comments, and up-vote to down-vote ratio. Approximately 64% of the samples have both text and image data (Nakamura et al., 2020). Samples of this dataset are annotated with 2-way, 3-way, and 6-way labels including true, satire/parody, misleading content, manipulated content, false connection, and imposter content. Examples of 6-way labels are demonstrated in Fig. 6. Additionally, Table 2 illustrates a comparison and evaluation of various methods’ performance on the Fakeddit dataset (Nakamura et al., 2020) [^15] The table is based on the work in (Nakamura et al., 2020).
@@
Table 2. Evaluation of classification accuracy on the Fakeddit dataset using various image/text embedders conducted by (Nakamura et al., 2020).
@@
(a) Caption: These squirrels fighting.
Class: True Content
(b) Caption: Incredible new images of Jupiter from NASA.
Class: Misleading Content
(c) Caption: Man binge-watches life passing him by. Class: Satire/Parody
(d) Caption: King Arthur and his noble steed.
Class: Manipulated Content
(e) Caption: Picture my sister left her corner after 30hrs, and had to go to first grade this year!
Class: False Connection
(f) Caption: “One Raid by an Enemy Bomber…” Leaflets dropped over US during Project Revere, an American Cold War propaganda test.
Class: Imposter Content
Figure 6. Examples of different classes in Fakeddit dataset (Nakamura et al., 2020).
NewsBag: comprises 200,000 real news and 15,000 fake articles. The real training articles have been collected from the Wall Street Journal and the fake ones from The Onion website [^16] https://www.theonion.com/, which publishes satirical content. However, the samples of the test set are collected from different websites, i.e., TheRealNews [^17] https://therealnews.com/ and ThePoke [^18] https://www.thepoke.co.uk/. The rationale behind using different sources of news for the training and test sets is to observe how well the models could be generalized to unseen data samples. The NewsBag dataset is highly imbalanced. Thus, to tackle this issue, NewsBag ++ is also released, which is the augmented training version of the NewsBag dataset and contains 200,000 real and 389,000 fake news articles. Another weakness of the NewsBag dataset is that it does not have any social context information such as spreader information, sharing trends, and reactions such as user comments and engagements (Jindal et al., 2020).
MM-COVID: [^19] https://github.com/bigheiniu/MM-COVID is a multi-lingual and multi-dimensional COVID-19 fake news data repository. This dataset comprises 3,981 fake news and 7,192 trustworthy information in 6 different languages, i.e., English, Spanish, Portuguese, Hindi, French, and Italian. MM-COVID consists of visual, textual, and social context information, e.g., users and networks information (Li et al., 2020a). This dataset is annotated is by Snopes202020www.snopes.com and Poynter212121www.poynter.org/coronavirusfactsalliance/ crowdsource domains where experts and journalists evaluate and fact-check news content and annotate contents as either fake or real. While Snopes is an independent publication that mainly contains English content, Poynter is an international fact-checking network (IFCN) which unites 96 different fact-checking agencies such as PolitiFact [^22] https://www.politifact.com/ in 40 languages.
ReCOVery: [^23] https://github.com/apurvamulay/ReCOVery contains 2,029 news articles that have been shared on social media, most of which (2,017 samples) have both textual and visual information for multi-modal studies. ReCOVery is imbalanced in news class, i.e., the proportion of real vs. fake articles is around 2:1. The number of users who spread real news (78,659) and users sharing fake articles (17,323) is greater than the total number of users included in the dataset (93,761). In this dataset, the assumption is that users can engage in spreading both real and fake news articles. Samples of this dataset are annotated by two fact-checking resources: NewsGuard [^24] https://www.newsguardtech.com/ and Media Bias/Fact Check (MBFC) [^25] https://mediabiasfactcheck.com/, which is a website that evaluates factual accuracy and political bias of news media. MBFC labels each news media as one of six factual-accuracy levels based on the fact-checking results of the previously published news articles. Samples of ReCOVery are collected from 60 news domains, from which 22 are the sources of reliable news articles (e.g., National Public Radio [^26] https://www.npr.org/ and Reuters [^27] https://www.reuters.com) and the remaining 38 are sources to collect unreliable news articles (e.g., Human Are Free [^28] http://humansarefree.com/ and Natural News [^29] https://www.naturalnews.com/(Zhou et al., 2020a).
CoAID: [^30] https://github.com/cuilimeng/CoAID Covid-19 heAlthcare mIsinformation Dataset or CoAID is a diverse COVID-19 healthcare misinformation dataset, including fake news on websites and social platforms, along with users’ social engagement about the news. It includes 5,216 news articles, 296,752 related user engagements, 926 social platform posts about COVID-19, and ground truth labels. The publishing dates of the collected information range from December 1, 2019, to September 1, 2020. In total, 204 fake news articles, 3,565 true news articles, 28 fake claims, and 454 true claims are collected. Real news articles are crawled from 9 reliable media outlets that have been cross-checked as reliable, e.g., National Institutes of Health (NIH) [^31] https://www.nih.gov/news-events/news-releases and CDC [^32] https://www.cdc.gov/coronavirus/2019-ncov/whats-new-all.html. Fake news is retrieved from several fact-checking websites, such as PolitiFact and Health Feedback [^33] https://healthfeedback.org/ (Cui and Lee, 2020).
MMCoVaR: is a Multi-modal COVID-19 Vaccine Focused Data Repository (MMCoVaR). Articles in this dataset are annotated using two news website source checking methods, and the tweets are fact-checked based on a stance detection approach. MMCoVaR comprises 2,593 articles issued by 80 publishers and shared between 02/16/2020 and 05/08/2021, and 24,184 Twitter posts collected between 04/17/2021 and 05/08/2021. Samples of this dataset are annotated by Media Bias Chart and Media Bias/Fact Check (MBFC) and classified into two levels of credibility: reliable and unreliable. Thus, articles are labeled as either credible or unreliable, and tweets are annotated as reliable, inconclusive, or unreliable (Chen et al., 2021). It is worth mentioning that textual, visual, and social context information are available for the news articles.
N24News: [^34] https://github.com/billywzh717/N24News is a multi-modal dataset extracted from the New York Times articles published from 2010 to 2020. Each news article belongs to one of 24 different categories, e.g., science, arts, etc. The dataset comprises up to 3,000 samples of real news for each category. In total, 60,000 news articles are collected. Each article sample contains a category tag, headline, abstract, article body, image, and corresponding image caption. This dataset is randomly split into training/validation/testing sets in the ratio of 8:1:1 (Wang et al., 2022b). The main weakness of this dataset is that it does not have any fake samples, and all of the real samples are collected from a single source, i.e., The New York Times.
@@
MuMiN: [^35] https://github.com/MuMiN-dataset/mumin-build Large-Scale Multilingual Multi-modal Fact-Checked Misinformation Social Network Dataset (MuMin) comprises 21 million tweets belonging to 26 thousand Twitter threads, each of which has been linked to 13 thousand fact-checked claims in 41 different languages. MuMiN is available in three versions: large, medium, and small, with the largest one consisting of 10,920 articles and 6,573 images. In this dataset, if the claim is “mostly true,” it is labeled as factual. When the claim is deemed “half true” or “half false,” it is labeled as misinformation, with the justification that a statement containing a significant part of false information should be considered misleading content. When there is no clear verdict, the verdict is labeled as other (Nielsen and McConville, 2022).
@@
A summary and side-by-side comparison of the previously mentioned datasets are shown in Table 3. As illustrated in Fig. 7, most of these datasets are small, annotated with binary labels, sourced from limited platforms like Twitter, and contain only a few modalities, namely text and image.
@@
N24News (Wang et al., 2022b) 60,000 24 image,text New York Times All samples are real from 24 different categories
@@
Figure 7. Number of news articles by dataset.
5. CHALLENGES IN MULTI-MODAL MISINFORMATION DETECTION
Recent studies on multi-modal learning have made significant contributions to the field of multi-modal fake news detection. However, there are still weaknesses and shortcomings, and recognizing them opens the door to new opportunities not only in fake news detection but also in the multi-modal field in general. In this section, we discuss challenges, shortcomings, and opportunities in multi-modal fake news detection. We provide non-exhaustive lists of challenges and shortcomings for each direction of multi-modal misinformation research.
5.1. Data study Challenges